OpenStudy (anonymous):

Help please :) Find the inverse of the following function: 5/2x -15

OpenStudy (anonymous):

the root is x = 6. @shadly

OpenStudy (anonymous):

.....What?? I don't need a root... :/

OpenStudy (anonymous):

What are the choices available for your question?

OpenStudy (anonymous):

2x-30/5, 2x+30/5, 2/5x+15, or 2/5x-15

OpenStudy (anonymous):

I have no idea... @JA1

OpenStudy (anonymous):

...Thanks .-.

OpenStudy (ja1):

Fractions type algebra is not my suit @melody16 @tcarroll010

OpenStudy (anonymous):

You'll get help in just a second please be patient. :) @shadly

OpenStudy (johnweldon1993):

The inverse of \[y = \frac{ 5 }{ 2 }x - 15\] first thing you do is switch the 'x' and the 'y' \[x = \frac{ 5 }{ 2 }y - 15\] Now solve for 'y' again...can you do that?

OpenStudy (anonymous):

I'm not really sure how which is why I asked the question... I know you have to switxh x and y but im kinda stuck

OpenStudy (anonymous):

switch*

OpenStudy (johnweldon1993):

Okay....well as shown above...I have switched them....but now to solve for y again...we need to isolate it \[x = \frac{ 5 }{ 2 }y - 15\] Lets start by adding 15 to both sides....now we'll have \[x + 15 = \frac{ 5 }{ 2 }y\] Now...we want to get rid of that fraction next to the 'y'....so multiply both sides on the equation by 2/5 (the reciprocal of 5/2) \[\frac{ 2x + 30 }{ 5 } = y\] So the inverse of your equation...is \[y = \frac{ 2x + 30 }{ 5 }\] or written as \[y = \frac{ 2 }{ 5 }x + 6\]

OpenStudy (anonymous):

ooooh okay... Thank you

OpenStudy (johnweldon1993):

No problem! Now did that make sense?

OpenStudy (anonymous):

Yes, a lot XD

OpenStudy (johnweldon1993):

Awesome!